# Dual Consolidation for Pre-Trained Model-Based Domain-Incremental Learning

### v1

@@ 4.2 Classifier Consolidation
In Eq. 5, we design the feature consolidation process, which aggregates multiple fine-tuned models by accumulating task vectors. However, a fatal problem still exists, i.e., there is a mismatch between the classifiers and the consolidated embeddings. Since the classifiers are optimized to match the embeddings to the corresponding class, the matching degree drastically decays as the backbone is replaced with another one. Recalling the background information in Section 3.1 that we utilize an independent classifier for each domain, we need to design the extra classifier consolidation process to align the classifiers to the embedding space. In each training stage, we denote the classifier for previous domains as ‘old classifier’ ($W_{o}$) and the classifier for the current domain as ‘new classifier’ ($W_{n}$). We then discuss how to align them with the consolidated features.
New Classifier Retraining: In each training stage, we have the training data $\mathcal{D}^{b}$ in hand. Consequently, it is intuitive to coordinate the new classifier with the consolidated features via:
@@
where the consolidated feature $\bar{\phi}_{i}^{m}(\cdot)$ is frozen to resist further mismatch, and we can align the new classifier ${W_{n}}$ to match the features with Eq. 8.
@@
Old Classifier Transport: Eq. 8 aligns the new classifier to the latest embedding space. However, a fatal problem still exists, i.e., the old classifier is also incompatible with the merged embedding space. Consequently, the previous knowledge of old domains shall be forgotten in the incremental learning process. If we have plenty of training instances of previous domains, a similar calibration process can be done as in Eq. 8. However, due to the exemplar-free restrictions, we cannot save any previous instances in the memory. Hence, we need to find another way to estimate the relative calibration weight for the old classifier. For example, if we have the retrained classifier to classify a ‘lion’ in the clip art style, it can be mostly reused to classify the lion in the real photo style. We call such a relationship ‘semantic information,’ and the goal is to utilize such information to assist old classifier alignment. We denote the calibrated classifier using semantic information as $\hat{W}_{o}=\mathcal{T}(W_{n},\textbf{S})$, i.e., the estimated classifier is obtained by transforming the new classifier using semantic information S. Hence, there remain two core problems to be solved: 1) How to define the transformation function $\mathcal{T}$? 2) How to define the semantic information S?
@@
Implementing $\mathcal{T}$ via Optimal Transport: The function $\mathcal{T}$ encodes the correlation between the set of features among two tasks. Considering that the weight matrix of a linear layer reveals the relative relationship among feature-class pairs and the final logit is the aggregation of all features, we can design the recombination of exiting classifiers with a linear mapping $T\in\mathbb{R}^{\beta\times\gamma}$. $T$ encodes the cross-task correlation between the current and the previous domains, where larger values in $T$ denote higher class-wise similarity. Since the decision is made by matching the classifier with corresponding features, we can utilize and recombine the weights of similar classes in the current domain to obtain those of previous domains. For example, important features that discriminate lions in the clip art style should be assigned higher coefficients to help classify lions in the photo style and vice versa.
@@
We denote $\bm{\mu}_{1}\in\Delta_{\beta}$ and $\bm{\mu}_{2}\in\Delta_{\gamma}$, where $\Delta_{d}=\left\{\bm{\mu}:\bm{\mu}\in\mathbb{R}_{+}^{d},\bm{\mu}^{\top}\mathbf{1}=1\right\}$ is the $d$-dimensional simplex. $\bm{\mu}_{1}$ and $\bm{\mu}_{2}$ denote the importance of each class across domains, and we set them as uniform distribution. To map a set of classes to another between old and new domains, a cost matrix $Q\in\mathbb{R}_{+}^{\beta\times\gamma}$ is further introduced to guide the transition. The larger weight of $Q_{i,j}$ indicates we need to pay more cost when reusing the classifier of $i$-th class to assist the $j$-th class. Consequently, the matrix $T$ can be formulated as the coupling of two distributions, aiming to connect classes from different domains at the lowest transportation cost. Hence, $T$ can be obtained via minimizing:
@@
which is the Kantorovich formulation of optimal transport (OT) Villani (2008); Peyré et al. (2019). Optimizing Eq. 9 enables us to show how to move the probability mass across domains with low cost. Since the target is to reuse the retrained classifier $W_{n}$ to adjust previous classifier $W_{o}$, we set $\bm{\mu}_{1}\in\Delta_{|Y|}$ and $\bm{\mu}_{2}\in\Delta_{|Y_{o}|}$ with uniform class marginals, where $|Y_{o}|$ denotes the number of classes in $W_{o}$. Solving Eq. 9 produces the alignment between different domains, and we can apply $T$ to the new classifier to obtain the estimated old classifier, i.e., $\hat{W}_{o}=W_{n}T$.
Defining the Transportation Cost with Semantic Information: Solving Eq. 9 requires a proper definition of the cross-domain cost, i.e., $Q$. The higher cost indicates it is less effective to transport the classifier to the target class and vice versa. To this end, we design a simple yet effective way to measure such class-wise information. Specifically, before the training of each stage, we utilize the pre-trained backbone $\phi_{0}$ to extract class centers in its embedding space via Eq. 6, i.e., $[{\bm{c}}^{0}_{1},{\bm{c}}^{0}_{2},\cdots,{\bm{c}}^{0}_{|{Y}|}]$. These class centers indicate the most representative embedding of the corresponding classes. Since $\phi_{0}$ is optimized with extensive datasets, we rely on its generalizability to reflect task-wise similarity. If two classes are similar, their embeddings should also be situated near each other. Consequently, we calculate the Euclidean distance between class centers as the transportation cost, i.e., $Q_{i,j}=\left\|{\bm{c}}^{0}_{i}-{\bm{c}}^{0}_{j}\right\|_{2}^{2}$. Here classes $i$ and $j$ are from different domains. With the pair-wise transportation cost, we are able to optimize Eq. 9 for the transportation plan $T$. Afterward, we utilize the interpolation between old classifiers and the transported classifier for consolidation:
@@
Effect of classifier consolidation: We visualize the classifier consolidation process in Figure 1 (bottom). The classifier consolidation process takes two steps, i.e., new classifier retraining and old classifier transport. The first step is designed to align the new classifiers with the unified embedding, while the second step aims to recombine the old classifiers with the calibrated classifier. In this way, we obtain a compatible classifier with the consolidated features, thus favoring the recognition in the unified embedding space.

### v2

@@ 4.2 Classifier Consolidation
In Eq. (7), we design the representation consolidation process, which aggregates multiple fine-tuned models by accumulating task vectors. However, a fatal problem still exists, i.e., there is a mismatch between the classifiers and the consolidated embeddings. Since the classifiers are optimized to align the embeddings with the corresponding class, the matching degree can decrease drastically when the embedding space is totally updated. Recalling the background information in Section 3.1 that we utilize an independent classifier for each domain, we need to design an extra classifier consolidation process to align the classifiers with the updated embedding space. In each training stage, we denote the classifier for previous domains as ‘old classifier’ ($W_{o}$) and the classifier for the current domain as ‘new classifier’ ($W_{n}$). We then discuss how to align them with the consolidated features.
New Classifier Retraining: In each training stage, we have the training data $\mathcal{D}^{b}$ in hand. Hence, it is intuitive to coordinate the new classifier with the consolidated features via:
@@
where the consolidated feature $\bar{\phi}_{i}^{m}(\cdot)$ is frozen to resist further mismatch, and we can adjust the new classifier ${W_{n}}$ to match the features using Eq. (8).
@@
Old Classifier Transport: Though Eq. (8) aligns the new classifier to the latest embedding space, another issue remains, i.e., the old classifier is also incompatible with the merged embedding space. Consequently, the previous knowledge of old domains shall be forgotten in the incremental learning process. If we have plenty of training instances of previous domains, a similar calibration process can be done as in Eq. (8). However, due to the exemplar-free restrictions, we cannot save any previous instances in the memory. Hence, we need to find another way to estimate the relative calibration weight for the old classifier. For example, if we have the retrained classifier to classify a ‘lion’ in the clip art style, it can be mostly reused to classify the lion in the real photo style. We call such a relationship ‘semantic information,’ and the goal is to utilize such information to assist old classifier alignment. We denote the calibrated classifier using semantic information as $\hat{W}_{o}=\mathcal{T}(W_{n},\textbf{S})$, i.e., the estimated classifier is obtained by transforming the new classifier using semantic information S. Hence, there remain two core problems to be solved: 1) How to define the transformation function $\mathcal{T}$? 2) How to define the semantic information S?
@@
Implementing $\mathcal{T}$ via Optimal Transport: The function $\mathcal{T}$ encodes the correlation between the set of features among two tasks. Considering that the weight matrix of a linear layer reveals the relative relationship among feature-class pairs and the final logit is the aggregation of all features, we can design the recombination of existing classifiers with a linear mapping $T\in\mathbb{R}^{\beta\times\gamma}$. $T$ encodes the cross-task correlation between the current and the previous domains, where larger values in $T$ denote higher class-wise similarity. Since the decision is made by matching the classifier with corresponding features, we can utilize and recombine the weights of similar classes in the current domain to obtain those of previous domains. For example, important features that discriminate lions in the clip art style should be assigned higher coefficients to help classify lions in the photo style and vice versa.
@@
We denote $\bm{\mu}_{1}\in\Delta_{\beta}$ and $\bm{\mu}_{2}\in\Delta_{\gamma}$, where $\Delta_{d}=\left\{\bm{\mu}:\bm{\mu}\in\mathbb{R}_{+}^{d},\bm{\mu}^{\top}\mathbf{1}=1\right\}$ is the $d$-dimensional simplex. $\bm{\mu}_{1}$ and $\bm{\mu}_{2}$ denote the importance of each class across domains, and we set them as uniform distributions. To map the distribution of $\beta$ new classes to that of $\gamma$ classes from old domains, a cost matrix $Q\in\mathbb{R}_{+}^{\beta\times\gamma}$ is further introduced to guide the transition. The larger weight of $Q_{i,j}$ indicates we need to pay more cost when reusing the classifier of $i$-th class to assist the $j$-th class. Consequently, the transport matrix $T$ can be formulated as the coupling of two distributions, aiming to connect classes from different domains at the lowest transportation cost. Hence, $T$ can be obtained via minimizing:
@@
which is the Kantorovich formulation of optimal transport (OT) [64, 47]. Optimizing Eq. (9) enables us to show how to move the probability mass across domains with low cost. Since the target is to reuse the retrained classifier $W_{n}$ to adjust previous classifier $W_{o}$, we set $\bm{\mu}_{1}\in\Delta_{|Y|}$ and $\bm{\mu}_{2}\in\Delta_{|Y_{o}|}$ with uniform class marginals, where $|Y_{o}|$ denotes the number of classes in $W_{o}$. Solving Eq. (9) establishes the mapping relationship between different domains, and we can apply $T$ to the new classifier to obtain the estimated old classifier, i.e., $\hat{W}_{o}=W_{n}T$.
Defining the Transportation Cost with Semantic Information: Solving Eq. (9) requires a proper definition of the cross-domain cost, i.e., $Q$. The higher cost indicates it is less effective to transport the classifier to the target class and vice versa. To this end, we design a simple yet effective way to measure such class-wise information. Specifically, before the training of each stage, we utilize the pre-trained backbone $\phi_{0}$ to extract class centers in its embedding space via Eq. (5), i.e., $[{\bm{c}}^{0}_{1},{\bm{c}}^{0}_{2},\cdots,{\bm{c}}^{0}_{|{Y}|}]$. These class centers indicate the most representative embedding of the corresponding classes. Since $\phi_{0}$ is optimized with extensive datasets, we rely on its generalizability to reflect task-wise similarity. If two classes are similar, their embeddings should also be situated near each other. Consequently, we calculate the Euclidean distance between class centers as the transportation cost, i.e., $Q_{i,j}=\left\|{\bm{c}}^{0}_{i}-{\bm{c}}^{0}_{j}\right\|_{2}^{2}$. Here, classes $i$ and $j$ are from different domains. With the pair-wise transportation cost, we are able to optimize Eq. (9) for the transportation plan $T$. Afterward, we utilize the interpolation between old classifiers and the transported classifier for consolidation:
@@
Effect of classifier consolidation: We visualize the classifier consolidation process in Figure 1 (bottom). The classifier consolidation process takes two steps, i.e., new classifier retraining and old classifier transport. The first step is designed to align the new classifiers with the unified embedding, while the second step aims to recombine the old classifiers with the calibrated classifier. In this way, we obtain a unified classifier compatible with the consolidated features, thus favoring the recognition in the unified embedding space.
(a) CDDB ViT-B/16 IN1K
(b) CORe50 ViT-B/16 IN1K
(c) DomainNet ViT-B/16 IN1K
Method Office-Home DomainNet CORe50 CDDB
Replay [51] 84.23$\pm$2.31 83.75$\pm$0.68 64.78$\pm$2.98 61.16$\pm$1.19 85.56$\pm$0.38 92.21$\pm$0.63 66.91$\pm$18.0 63.21$\pm$11.6
iCaRL [52] 81.66$\pm$2.43 81.11$\pm$1.23 59.89$\pm$2.86 57.46$\pm$2.31 74.43$\pm$3.18 79.86$\pm$2.96 68.43$\pm$18.7 70.50$\pm$16.5
MEMO [85] 71.18$\pm$2.76 63.09$\pm$1.80 61.92$\pm$5.39 58.41$\pm$3.20 64.80$\pm$3.16 68.24$\pm$2.41 60.87$\pm$13.4 58.09$\pm$11.7
SimpleCIL [88] 75.69$\pm$5.03 75.72$\pm$0.00 42.95$\pm$4.84 44.08$\pm$0.00 70.92$\pm$0.74 74.80$\pm$0.00 60.80$\pm$4.49 63.40$\pm$0.00
L2P [70] 79.72$\pm$4.19 80.03$\pm$1.29 50.45$\pm$4.10 48.72$\pm$2.83 83.57$\pm$0.35 87.87$\pm$0.51 67.33$\pm$5.98 64.45$\pm$5.83
DualPrompt [69] 80.20$\pm$3.81 80.85$\pm$0.14 52.28$\pm$3.35 50.46$\pm$3.17 84.53$\pm$0.89 87.27$\pm$1.06 68.33$\pm$7.52 71.41$\pm$1.24
CODA-Prompt [58] 84.70$\pm$2.94 85.07$\pm$0.34 59.85$\pm$4.49 59.99$\pm$0.88 87.92$\pm$0.41 91.57$\pm$0.69 69.19$\pm$6.11 74.18$\pm$1.33
EASE [86] 81.16$\pm$3.52 76.33$\pm$2.16 50.50$\pm$2.27 43.72$\pm$1.70 86.30$\pm$0.04 87.02$\pm$1.21 67.78$\pm$2.44 64.96$\pm$8.36
RanPAC [40] 82.30$\pm$3.34 82.28$\pm$0.00 55.20$\pm$3.93 54.80$\pm$0.36 79.16$\pm$0.55 81.38$\pm$0.12 78.92$\pm$4.85 80.48$\pm$0.68
S-iPrompt [68] 81.50$\pm$3.17 80.51$\pm$0.21 61.16$\pm$4.57 60.46$\pm$0.90 81.95$\pm$0.57 83.38$\pm$0.70 68.51$\pm$7.20 72.76$\pm$0.43
@@ 4.3 Summary of Duct
In Duct, facing a new training task, we extract the class centers with $\phi_{0}$. Afterward, we optimize the model and separately consolidate the features and classifiers. The consolidated model ($[W_{o}^{m};W_{n}]^{\top}\phi^{m}_{i}$) is then utilized for testing. We utilize Sinkhorn’s algorithm [57] to solve the OT problem in Eq. (9). We summarize the training steps with pseudo code in the supplementary material.
Discussions on memory cost: During the training process, we need to keep two models in the memory, i.e., the historical consolidated backbone $\phi^{m}_{i-1}$ and the online backbone $\phi_{i}$ for learning the current task. However, after the learning process of each domain, the current running model is merged into $\phi^{m}_{i}$ (Eq. (7)) without the need to be stored in memory. During inference, we utilize the consolidated backbone $\phi^{m}_{i}$ and classifiers $[W_{o}^{m};W_{n}]$ for classification, which is the same as a single backbone. Consequently, Duct can be fairly compared with others.
5 Experiment
(a) Multiple trails
(b) Forgetting measure
(c) Parameter robustness
5.1 Experimental Setup
Dataset: Following the benchmark settings [69, 70, 68, 58] in PTM-based domain-incremental learning, we evaluate the performance on Office-Home [63], DomainNet [46], CORe50 [36], and CDDB-Hard [30]. Specifically, Office-Home has four domains, DomainNet has six domains, CORe50 has 11 domains, and CDDB-Hard has five domains. We consider five task orders to shuffle these domains in the DIL setting for a holistic evaluation and report the details in the supplementary material.
Comparison methods: In the comparison, we consider two types of methods, including exemplar-based methods, i.e., Replay [51], iCaRL [52], MEMO [85], and SOTA exemplar-free DIL methods, i.e., SimpleCIL [88], L2P [70], DualPrompt [69], CODA-Prompt [58], EASE [86], RanPAC [40], and S-iPrompt [68]. We utilize the same pre-trained backbone for all compared methods.
Implementation details: We deploy the experiments using PyTorch [45] and Pilot [59] on NVIDIA 4090. Following [70, 86], we consider two typical pre-trained weights, i.e., ViT-B/16-IN21K and ViT-B/16-IN1K. Both are pre-trained with ImageNet21K [54], while the latter is further finetuned on ImageNet1K. We optimize Duct using SGD optimizer with a batch size of 128 for 15 epochs. The learning rate is set to $0.001$. We select 10 exemplars per class for exemplar-based methods using herding [71] algorithm. In Duct, we set the consolidation parameter $\alpha_{\phi}=0.5,\alpha_{w}=0.5$. The code will be made publicly available upon acceptance.
Performance Measure: Following [70, 68], we denote the accuracy among all seen domains after learning the $b$-th task as $\mathcal{A}_{b}$, we mainly consider $\mathcal{A}_{B}$ (the performance after learning the last stage) and $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$ (the average performance among all incremental stages) for comparison. We also consider the forgetting measure [13] to measure the relative forgetting degree in DIL.
5.2 Benchmark Comparison
We report the results on benchmark datasets in Figure 2 and Table 1, with the remaining results provided in the supplementary material. As we can infer from these results, Duct consistently outperforms other compared methods by $\bf 1\sim 7$in the final accuracy. It must be noted that the differences of the starting point are due to the different tuning techniques, which cannot be directly aligned since the number of trainable parameters are different in those methods. Specifically, sequentially finetuning the model suffers from catastrophic forgetting, even starting with the pre-trained model. To compensate for the forgetting phenomena, we observe that several exemplar-based methods (Replay, iCaRL, and MEMO) show stable improvements to the baseline. However, since some of these works are designed for the class-incremental learning scenario, the expansion or distillation target may not be optimal for the current setting. We then compare Duct to the methods specially designed for PTMs and find prompt-based methods (L2P, DualPrompt, CODA-Prompt, and S-iPrompt) yield inferior performance to ours. We also observe that Duct is compatible with various pre-trained weights and shows stable improvements with ViT-B/16 IN1K and IN21K (see supplementary).
Besides, we also report the incremental performance among five task orders (average and standard deviation) in Table 1 and Figure 3(a). As we can infer from the table, Duct works robustly across different task orders, showing stable improvements against other state-of-the-art methods.
Variations $\bar{\mathcal{A}}$ ${\mathcal{A}}_{B}$
Baseline 66.56 63.40
Variation 1 80.36 74.17
Variation 2 81.41 76.82
Variation 3 85.42 80.31
Duct 87.74 82.35
Table 2: Ablation study on differnt modules in Duct.
Figure 4: Before Duct.
Figure 5: After Duct.
Figure 6: Visualizations of the embedding space via t-SNE [62] on DomainNet. We visualize two incremental stages and utilize dots to represent the first domain and triangles for the second domain. The shadow region denotes the decision boundaries.
5.3 Further Analysis
Forgetting Measure: Figure 2 and Table 1 mainly focus on the accuracy measure, which utilizes all seen domains to measure the relative performance. Apart from the accuracy measure, we also follow [68, 70] to utilize the forgetting measure, which captures the stability of existing knowledge among previous domains. As shown in Figure 3(b), we report the forgetting measure among all compared methods on the CDDB dataset. As we can infer from the figure, several methods suffer from severe forgetting, e.g., Finetune and MEMO, indicating that they are unsuitable for the domain-incremental learning scenario. We also observe that prompt-based methods freeze the backbone to resist feature drift, which tackles the forgetting phenomena. However, Duct still shows the least forgetting (i.e., 0.12) among all compared methods, indicating its strong performance.
Parameter Robustness: There are two major consolidation steps in Duct, i.e., the representation consolidation in Eq. (7) and the classifier consolidation in Eq. (10). These consolidation steps involve the hyper-parameters for merging a set of models/classifiers, i.e., the backbone merge ratio $\alpha_{\phi}$ and the classifier merge ratio $\alpha_{W}$. In this section, we conduct ablations on the choice of these parameters to investigate the robustness of Duct to their changes. Specifically, we choose $\alpha_{\phi}$ and $\alpha_{W}$ among $\{0.1,0.25,0.5,0.75,1.0\}$, resulting in 25 parameter combinations. We conduct experiments with these parameter combinations on the Office-Home dataset and present the average performance in Figure 3(c). As we can infer from the figure, Duct is generally robust to parameter changes. Besides, since the merge parameters are designed to trade-off between old and new knowledge, we find either a small value (e.g., 0.1) or a large value (e.g., 1.0) works poorly. The poor performance of these marginal parameters also indicates the importance of the merging process since a small value or large value equals ignoring part of the important information during consolidation. By contrast, both of them prefer a mild value, i.e., choosing them around {0.25, 0.5} shows better performance. Hence, we suggest setting $\alpha_{\phi}=0.5,\alpha_{w}=0.5$ as the default setting.
Compositional Components: In this section, we conduct experiments to investigate the importance of each module in Duct on CDDB dataset. As shown in Table 2, ‘Baseline’ denotes directly freezing the embedding and extracting class centers as the classifier, which shows poor performance due to the domain gap to the downstream domains. We then utilize Eq. (4) to sequentially merge the backbone to consolidate the representations and denote it as Variation 1. It shows that adding the representation consolidation process drastically improves the performance, indicating the superiority of such unified representation in DIL. However, when switching Eq. (4) to Eq. (7), we find Variation 2 further improves the performance, indicating that task similarity is helpful in obtaining a universal embedding space. Afterward, we combine it with the classifier retraining process in Eq. (8), which is denoted as Variation 3. It shows that the mismatch between consolidated features and the classifiers weakens the performance, and aligning the new classifiers to the embedding helps DIL. When comparing Duct to Variation 3, we find the old classifier transport is also vital to recover former knowledge and resist forgetting, which improves the final accuracy by 2%. Ablation studies verify the efficacy of different modules in Duct.
Visualizations: In this section, we visualize the embedding space to show the effectiveness of Duct on the DomainNet dataset. We consider a two-stage domain-incremental learning scenario, each containing five classes, and utilize t-SNE [62] to visualize the representation space before and after Duct. We use dots to represent the classes in the first domain and triangles for classes in the second domain. We utilize the shadow regions to denote the decision boundary obtained by the classifier weights to represent different classes. As shown in Figure 4, although the embedding space before Duct shows competitive performance, there are two major flaws: 1) There exists the confusion region between yellow and red dots, indicating that previous classes are partially forgotten as data evolves. 2) The embedding space of the same class is still located in different regions, e.g., the purple and green areas, which is not ideal for domain-incremental learning. However, when applying Duct to consolidate the features and classifiers, the above problems are addressed in Figure 5. Specifically, since the consolidated embedding suits all tasks, the forgetting of previous domains is alleviated. Besides, the unified embedding space adequately clusters the same class of different domains together, favoring the final inference.
6 Conclusion
Domain-incremental learning is a desired ability for real-world learning systems to obtain new knowledge. In this paper, we propose dual consolidation (Duct) to achieve this goal. Since the forgetting in DIL occurs in two aspects, i.e., the representation and the classifier, we separately design the consolidation technique to handle them. Specifically, we consider continually merging the pre-trained model with domain-specific task vectors to achieve the unified embedding. To compensate for the mismatch between consolidated features and classifiers, we design the classifier consolidation process by introducing semantic-guided transport. Extensive experiments validate the effectiveness of Duct.
Limitations: Possible limitations include the utilization of PTMs since representation consolidation relies on generalizable initialization. Future work includes extending Duct to non-PTM scenarios.
Acknowledgments. This work is partially supported by NSFC (U23A20382, 62476123, 62376118, 62006112, 62250069), Fundamental Research Funds for the Central Universities (2024300373,14380021), CCF-Tencent Rhino-Bird Open Research Fund RAGR20240101, Collaborative Innovation Center of Novel Software Technology and Industrialization.
Appendix
In the main paper, we present a method to prevent forgetting in domain-incremental learning through representation and classifier consolidation. The supplementary material provides additional details on the algorithm and experimental results mentioned in the main paper, along with extra empirical evaluations and discussions. The organization of the supplementary material is as follows:
  • Section I provides the pseudocode of Duct.
  • Section II presents detailed experimental results, including the running time comparison, detailed performance on different task orders, and results with other pre-trained weights.
  • Section III introduces the details about the datasets adopted in the main paper, including the number of tasks and images and differently-ordered sequences of tasks.
  • Section IV introduces the compared methods adopted in the main paper.
  • Section V presents a series of additional experiments to further validate the robustness of Duct.
Appendix I Algorithm Pipeline
Following the notation presented in the main body, we summarize the training steps in Algorithm 1.
First of all, we prepare the dataset and load the pre-trained model weights, then calculate the mean representation for each incoming class, as shown in Lines 2$\sim$3. Next, we fine-tune the model and extract the task vector, which is then used to apply the model merging technique, as described in Line 5. After that, we keep the merged backbone frozen and retrain the classification layer to align it with the consolidated representation. Finally, we apply classifier transport to the classifiers to transfer semantic information, as outlined in Lines 7$\sim$8.